OpenStudy (anonymous):

the equation of line AB is y=5x+1. write an equation of a line parallel to line AB in slope-intercept form that contains point (4,5).

OpenStudy (texaschic101):

ok...look at your equation y = 5x + 1. Can you tell me what the slope is ?

OpenStudy (texaschic101):

I am in a good mood....slope is 5. Parallel line will have same slope. Now use y = mx + b slope(m) = 5 (4,5) x = 4 and y = 5 now we sub and solve for b 5 = 5(4) + b 5 = 20 + b 5 - 20 = b -15 = b so your parallel line will be : y = 5x - 15

OpenStudy (texaschic101):

understand ?

OpenStudy (anonymous):

yes thankyou (: